### v1

@@ Appendix A Omitted statements and proofs from Section 4
Given an economy $E^{T}$, a period $t$, and a matching through period $t$, $m^{t-1}$. The one-period economy induced by the collection $\{\varphi_{t}(m^{t-1},k):k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$ is the length-1 economy defined as follows:
@@
      For each agent $b\in\pazocal{B}_{t}(m^{t-1})$, $v_{\varphi_{t}}(a,b)\geq v_{\varphi_{t}}(b,b)$ only if $v(a,b)\geq\min\{V_{t}(b,\overline{m}):\overline{m}\in\varphi_{t}(m^{t-1},b)\}$. Moreover, for $a,a^{\prime}\in A_{1}$, $v_{\varphi_{t}}(a,b)\geq v_{\varphi_{t}}(a^{\prime},b)$ only if $v(a,b)\geq v(a^{\prime},b)$.
@@
$M^{\star}(\Phi)=\left\{m^{\star}\in\pazocal{M}_{T}:(\forall t\in\{1,\dots,T\})m^{\star}_{t}\in\mathrm{S}\left(E_{t,\varphi_{t}}(m^{\star^{t-1}})\right)\right\},$
@@
$m^{\star}\in\varphi_{t}(k,m^{\star^{t-1}}).$ (A.1)
@@
Finally, suppose $m^{\star}_{t}$ fails part 3 of 6 for some pair $(a,b)$ in $\pazocal{A}_{t}(m^{\star^{t-1}})\times\pazocal{B}_{t}(m^{\star^{t-1}})$. Then at least one of the members of the pair must be single under $m^{\star}_{t}$. Without loss of generality, assume $m^{\star}_{t}(a)=a$. Then, $u(a,b)>U_{t}(a,m^{\star})\geq\min\{U_{t}(a,\overline{m}):\overline{m}\in\varphi_{t}(a,m^{\star^{t-1}})\}$, where the second inequality follows from $m^{\star}\in\varphi_{t}(a,m^{\star^{t-1}})$. This chain of inequalities contradicts the definition of $m^{\star}_{t}$. Thus $(m_{s}^{\star})_{s=t}^{T}\in\Sigma_{T-t+1}^{\Phi}(E^{T}_{t}(m^{\star^{t-1}}))$ and the result follows. ∎
@@
In Section 4, we argued that the period-1 conjectures defined in Equation 4.1 lead to nonempty solutions for $E^{T}$ provided that the solution concept for the continuation economy, $\Sigma_{T-1}^{\Phi}$ is also nonempty. For completeness, we now build recursively the conjectures and the solution concept to complete the argument: A solution concept as in 6 built off of conjectures that rule out matchings that are not consistent with the solution concept in the continuation is always nonempty. In what follows, we keep the notation $\varphi^{*}$ for the conjectures and denote by $\Sigma^{\Phi^{*}}$ for the induced solution concept.
Fix an economy of length $T$. Recall that 1 implies that no matter what the conjectures are, the set of $\Phi$-solutions for an economy of length $1$ coincide with the set of stable matchings. Given this, fix a matching through period $T-1$, $m^{T-2}$. For $k\in\pazocal{A}_{T-1}\left(m^{T-2}\right)\cup\pazocal{B}_{T-1}\left(m^{T-2}\right)$, define $k$’s conjectures as follows:
@@
$\varphi^{*}_{t}(m^{t-1},k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1},k):(\overline{m}_{s})_{s=t+1}^{T}\in\Sigma^{\Phi^{*}}_{T-t}\left(E^{T}_{t}(\overline{m}^{t})\right)\right\},$
and let $\Sigma^{\Phi^{*}}_{T-t+1}\left(E^{T}_{t}(m^{t-1})\right)$ denote the set of solutions according to the family $\Phi^{*}_{\geq t}=\{\varphi^{*}_{s}((m^{t-1},\overline{m}_{t}^{s-1}),k):s\geq t,\overline{m}_{t}^{s-1},k\in\pazocal{A}_{s}(m^{t-1},m_{t}^{s-1})\cup\pazocal{B}_{s}(m^{t-1},m_{t}^{s-1})\}$.
@@
For all $T\in\mathbb{N}$, $\Sigma^{\Phi^{*}}_{T}$ is a nonempty valued correspondence.
A.1 then shows that existence issues for solution concepts in dynamic matching markets do not arise solely from the perfection requirement–i.e., that the continuation matching also ascribes by the solution concept–but from restrictions the conjectures place on what period-$t$ matchings are feasible.
@@
It follows that $m_{\emptyset}$ is the unique rational expectations solution in any two-period economy in which at least $b_{1},b_{2}$ move to $t=2$.
@@
Third, consider $E_{\setminus\{a_{3},b_{1}\}}^{2}$. We show that $b_{1}$ is always matched to $a_{4}$ in any rational expectations solution and hence $b_{1}$ cannot match with $a_{1}$ in a rational expectations solution for $E_{\setminus\{a_{3}\}}^{2}$. In $E_{\setminus\{a_{3},b_{1}\}}^{2}$, three possibilities exist for a matching that is a rational expectations solution: (i) no one matches in $t=1$, which would lead to $m_{\emptyset}$ (though note that $m_{\emptyset}$ is blocked by $(a_{1},b_{2})$ in $E_{\setminus\{a_{3},b_{1}\}}^{2}$), (ii) $a_{1}$ matches with $b_{2}$, and (iii) $a_{2}$ matches with $b_{2}$. Under (ii), note that the outcome of side-$B$ proposing deferred acceptance in $t=2$ among $a_{2},a_{3},a_{4},b_{1},b_{3},b_{4}$ is
@@
so that any rational expectations solution for $E_{\setminus\{a_{3},b_{1}\}}^{2}$ that matches $a_{1},b_{2}$ in $t=1$ (if any) gives $b_{1}$ at least the payoff of matching with $b_{4}$. Similarly, under (iii), note that the outcome of side-$B$ proposing deferred acceptance in $t=2$ among $a_{1},a_{3},a_{4},b_{1},b_{3},b_{4}$ is
@@
Fourth, consider $E_{\setminus\{a_{2}\}}$. We argue that $m_{a_{2}}$ is a rational expectations solution for that economy. To see this, we need to argue that $b_{1}$ cannot block. Consider then the economy $E_{\setminus\{a_{2},b_{1}\}}$ and note that $m_{b_{1}}$ is a rational expectations solution for the same reason it is a rational expectations solution in $E_{\setminus\{b_{1}\}}^{2}$. Thus, by blocking $m_{a_{2}}$, $b_{1}$ cannot expect to get more than the payoff of matching with $a_{1}$ in $t=2$. It follows that $m_{a_{2}}$ is a rational expectations solution for $E_{\setminus\{a_{2}\}}^{2}$.
@@ Appendix B Omitted proofs from Section 5
Assume then that we know the correspondence $\hat{\mathrm{DS}}_{t}$ is nonempty valued for $t\in\{1,\dots,T-1\}$. As it is clear from the main text, to establish the main result, it is enough to show that for any $E^{T}$, the period-1 conjectures $\{\varphi_{\hat{\mathrm{DS}},1}(\cdot,k):k\in A_{1}\cup B_{1}\}$ are nonempty and satisfy consistency.
Fix an economy $E^{T}$. For each $k\in A_{1}\cup B_{1}$, let $\varphi_{0}(\cdot,k)=\{\overline{m}\in\pazocal{M}_{T}(\cdot,k):(\overline{m}_{t})_{t=2}^{T}\in\hat{\mathrm{DS}}_{T-1}(E^{T}_{2}(\overline{m}_{1}))\}$ and recursively define for $n\geq 1$
$\varphi_{n}(\cdot,k)=\left\{\overline{m}\in\pazocal{M}_{T}(\cdot,k):\begin{array}[]{l}\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{n-1}}^{\overline{m}_{1}})\\ (\overline{m}_{t})_{t=2}^{T}|_{E_{2}^{T}(\overline{m}_{1})}\in\hat{\mathrm{DS}}_{T-1}(E_{2}^{T}(\overline{m}_{1}))\end{array}\right\}.$
We observe that for all $k\in A_{1}\cup B_{1}$, (i) $\varphi_{n+1}(\cdot,k)\subseteq\varphi_{n}(\cdot,k)$ because in each step we are increasing the continuation values of period-1 agents other than $k$, (ii) because everything is finite $\varphi_{\infty}(\cdot,k)=\lim_{n\rightarrow\infty}\varphi_{n}(\cdot,k)$ is well-defined, and (iii) because $\{\overline{m}\in\pazocal{M}_{T}(\cdot,k):(\overline{m}_{t})_{t=2}^{T}\in\hat{\mathrm{DS}}_{T-1}\left(E_{2}^{T}(\overline{m}_{1})\right)\}\subseteq\varphi_{n}(\cdot,k)$ for all $n$ and is nonempty, then $\varphi_{\infty}(\cdot,k)$ is nonempty.
@@
$\varphi_{\infty}(\cdot,k)=\left\{\overline{m}\in\pazocal{M}_{T}(\cdot,k):\begin{array}[]{l}\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{\infty}}^{\overline{m}_{1}})\\ (\overline{m}_{s})_{s=2}^{T}|_{E_{2}^{T}(\overline{m}_{1})}\in\hat{\mathrm{DS}}(E_{2}^{T}(\overline{m}_{1}))\end{array}\right\},$ (B.3)
@@
    If $\overline{N}$ exists such that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\overline{N}}(\cdot,k)=\varphi_{\overline{N}+1}(\cdot,k)$, then for all $N\geq\overline{N}$ and all $k\in A_{1}\cup B_{1}$, it follows that $\varphi_{N}(\cdot,k)=\varphi_{\overline{N}+1}(\cdot,k)$. This follows from the definition of $\varphi_{n}$.
@@
    Because the sets $\varphi_{n}(\cdot,k)$ are finite and so is the set of period-1 agents, $\overline{N}<\infty$ exists such that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\overline{N}}=\varphi_{\overline{N}+1}$. Toward a contradiction, suppose that for all $N\in\mathbb{N}$ $k_{N}\in A_{1}\cup B_{1}$ exists such that $\varphi_{N}(\cdot,k_{N})\neq\varphi_{N+1}(\cdot,k_{N})$. Since the set $A_{1}\cup B_{1}$ is finite, we can find an agent $k$ and a subsequence $(n^{k})$ such that $\varphi_{n^{k}}(\cdot,k)\neq\varphi_{n^{k}+1}(\cdot,k)$. Thus, $\varphi_{n^{k}+1}(\cdot,k)\subsetneq\varphi_{n^{k}}(\cdot,k)$. Since $\varphi_{n^{k}}(k)$ is finite, this means that eventually $\varphi_{n^{k}}(k)=\emptyset$, a contradiction. From now on, let $\overline{N}$ denote the smallest $n$ such that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\overline{N}}(\cdot,k)=\varphi_{\overline{N}+1}(\cdot,k)\equiv\varphi_{\infty}(\cdot,k)$.
@@
    Suppose $k\in A_{1}\cup B_{1}$ exists such that $\overline{m}\in\varphi_{\infty}(\cdot,k)$, but $\overline{m}$ is not an element of the set on the right-hand side of Equation B.3. Then, without loss of generality, $a\in A_{1}\cup B_{1}\setminus\{k\}$ exists such that $\overline{m}_{1}(a)\neq a$ and
    $\min_{\tilde{m}\in\varphi_{\infty}(a)}U_{1}(a,\tilde{m})>u(a,\overline{m}_{1}(a)).$
    This implies that $\overline{m}\notin\varphi_{\overline{N}+2}(k)$, contradicting that $\varphi_{\overline{N}+1}(k)=\varphi_{\infty}(k)$.
We conclude that for all $k\in A_{1}\cup B_{1}$, Equation B.3 holds and hence $\varphi_{\infty}=\varphi_{\hat{\mathrm{DS}},1}$. Thus, for all $k\in A_{1}\cup B_{1}$, $\varphi_{\hat{\mathrm{DS}}}(k,\cdot)\neq\emptyset$.
We now show that $\varphi_{\hat{\mathrm{DS}},1}$ satisfy CC. To this end, construct a candidate $m^{\star}$ as described in Section 4 using $\varphi_{\hat{\mathrm{DS}},1}$ and the correspondence $\hat{\mathrm{DS}}_{T-1}$, which by the inductive hypothesis is nonempty. Fix a period-1 agent, $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)=k$. Note that $m^{\star}$ satisfies the following: First, the continuation matching $(m^{\star}_{t})_{t=2}^{T}\in\hat{\mathrm{DS}}(E^{T}_{2}(m^{\star}_{1}))$. Second, for all $k^{\prime}\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k^{\prime})\neq k^{\prime}$, we have that $k^{\prime}$ is matched to someone they prefer to the worst matching in $\varphi_{\hat{\mathrm{DS}},1}(\cdot,k^{\prime})$. Finally, note that $m^{\star}_{1}$ has no blocking pairs $(a,b)$ such that $m^{\star}_{1}(a)\neq a,m^{\star}_{1}(b)\neq b$. Thus, $m^{\star}_{1}\in\mathrm{S}(E_{1,\varphi_{\hat{\mathrm{DS}}}}^{m^{\star}_{1}})$. Hence, $m^{\star}\in\varphi_{\hat{\mathrm{DS}},1}(\cdot,k)$ so that CC holds.

### v2

@@ Appendix A Omitted statements and proofs from Section 4
Given an economy $E^{T}$, a period $t$, and a matching through period $t-1$, $m^{t-1}$. The one-period economy induced by the collection $\{\varphi_{t}(m^{t-1},k):k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$ is the length-1 economy defined as follows:
@@
      For each agent $b\in\pazocal{B}_{t}(m^{t-1})$, $v_{\varphi_{t}}(a,b)\geq v_{\varphi_{t}}(b,b)$ only if $v(a,b)\geq\min\{V_{t}(b,\overline{m}):\overline{m}\in\varphi_{t}(m^{t-1},b)\}$. Moreover, for $a,a^{\prime}\in\pazocal{A}_{t}(m^{t-1})$, $v_{\varphi_{t}}(a,b)\geq v_{\varphi_{t}}(a^{\prime},b)$ only if $v(a,b)\geq v(a^{\prime},b)$.
@@
$M^{\star}(\Phi)=\left\{m^{\star}\in\pazocal{M}_{T}:(\forall t\in\{1,\dots,T\})m^{\star}_{t}|_{E_{t,\varphi_{t}}(m^{\star^{t-1}})}\in\mathrm{S}\left(E_{t,\varphi_{t}}(m^{\star^{t-1}})\right)\right\},$
@@
$m^{\star}\in\varphi_{t}(m^{\star^{t-1}},k).$ (A.1)
@@
Finally, suppose $m^{\star}_{t}$ fails part 3 of 6 for some pair $(a,b)$ in $\pazocal{A}_{t}(m^{\star^{t-1}})\times\pazocal{B}_{t}(m^{\star^{t-1}})$. Then at least one of the members of the pair must be single under $m^{\star}_{t}$. Without loss of generality, assume $m^{\star}_{t}(a)=a$. Then, $u(a,b)>U_{t}(a,m^{\star})\geq\min\{U_{t}(a,\overline{m}):\overline{m}\in\varphi_{t}(m^{\star^{t-1}},a)\}$, where the second inequality follows from $m^{\star}\in\varphi_{t}(m^{\star^{t-1}},a)$. This chain of inequalities contradicts the definition of $m^{\star}_{t}$. Thus $(m_{s}^{\star})_{s=t}^{T}\in\Sigma_{T-t+1}^{\Phi}(E^{T}_{t}(m^{\star^{t-1}}))$ and the result follows. ∎
@@
In Section 4, we argued that the period-1 conjectures defined in Equation 4.1 lead to nonempty solutions for $E^{T}$ provided that the solution concept for the continuation economy, $\Sigma_{T-1}^{\Phi}$ is also nonempty. For completeness, we now build recursively the conjectures and the solution concept to complete the argument: A solution concept as in 6 built off of conjectures that rule out matchings that are not consistent with the solution concept in the continuation is always nonempty. In what follows, we keep the notation $\varphi^{*}$ for the conjectures and denote by $\Sigma^{\Phi^{*}}$ the induced solution concept.
Fix an economy of length $T$. Recall that 1 implies that no matter what the conjectures are, the set of $\Phi$-solutions for an economy of length $1$ coincide with the set of stable matchings. Given this, fix a matching through period $T-2$, $m^{T-2}$. For $k\in\pazocal{A}_{T-1}\left(m^{T-2}\right)\cup\pazocal{B}_{T-1}\left(m^{T-2}\right)$, define $k$’s conjectures as follows:
@@
$\varphi^{*}_{t}(m^{t-1},k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1},k):(\overline{m}_{s})_{s=t+1}^{T}\in\Sigma^{\Phi^{*}}_{T-t}\left(E^{T}_{t+1}(\overline{m}^{t})\right)\right\},$
and let $\Sigma^{\Phi^{*}}_{T-t+1}\left(E^{T}_{t}(m^{t-1})\right)$ denote the set of solutions given the family $\Phi^{*}_{\geq t}=\{\varphi^{*}_{s}((m^{t-1},\overline{m}_{t}^{s-1}),k):s\geq t,\overline{m}_{t}^{s-1},k\in\pazocal{A}_{s}(m^{t-1},m_{t}^{s-1})\cup\pazocal{B}_{s}(m^{t-1},m_{t}^{s-1})\}$.
@@
For all $T\in\mathbb{N}$, $\Sigma^{\Phi^{*}}_{T}$ is a nonempty-valued correspondence.
A.1 then shows that existence issues for solution concepts in dynamic matching markets do not arise solely from the perfection requirement–i.e., that the continuation matching also abides by the solution concept–but from restrictions the conjectures place on what period-$t$ matchings are feasible.
@@
It follows that $m_{\emptyset}$ is the unique rational expectations solution in any two-period economy in which at least $b_{1},b_{2}$ are part of the $t=2$-arrivals.
@@
Third, consider $E_{\setminus\{a_{3},b_{1}\}}^{2}$. We show that $b_{1}$ is always matched to $a_{4}$ in any rational expectations solution and hence $b_{1}$ cannot match with $a_{1}$ in a rational expectations solution for $E_{\setminus\{a_{3}\}}^{2}$. In $E_{\setminus\{a_{3},b_{1}\}}^{2}$, three possibilities exist for a matching that is a rational expectations solution: (i) no one matches in $t=1$, which would lead to $m_{\emptyset}$ (though note that $m_{\emptyset}$ is blocked by $(a_{1},b_{2})$ in $E_{\setminus\{a_{3},b_{1}\}}^{2}$), (ii) $a_{1}$ matches with $b_{2}$, and (iii) $a_{2}$ matches with $b_{2}$. Under (ii), note that the outcome of side-$A$ proposing deferred acceptance in $t=2$ among $a_{2},a_{3},a_{4},b_{1},b_{3},b_{4}$ is
@@
so that any rational expectations solution for $E_{\setminus\{a_{3},b_{1}\}}^{2}$ that matches $a_{1},b_{2}$ in $t=1$ (if any) gives $b_{1}$ at least the payoff of matching with $a_{4}$. Similarly, under (iii), note that the outcome of side-$A$ proposing deferred acceptance in $t=2$ among $a_{1},a_{3},a_{4},b_{1},b_{3},b_{4}$ is
@@
Fourth, consider $E_{\setminus\{a_{2}\}}^{2}$. We argue that $m_{a_{2}}$ is a rational expectations solution for that economy. To see this, we need to argue that $b_{1}$ cannot block. Consider then the economy $E_{\setminus\{a_{2},b_{1}\}}^{2}$ and note that $m_{b_{1}}$ is a rational expectations solution for the same reason it is a rational expectations solution in $E_{\setminus\{b_{1}\}}^{2}$. Thus, by blocking $m_{a_{2}}$, $b_{1}$ cannot expect to get more than the payoff of matching with $a_{1}$ in $t=2$. It follows that $m_{a_{2}}$ is a rational expectations solution for $E_{\setminus\{a_{2}\}}^{2}$.
@@ Appendix B Omitted proofs from Section 5
Assume then that we know the correspondence $\hat{\mathrm{DS}}_{t}$ is nonempty valued for $t\in\{1,\dots,T-1\}$. As it is clear from the main text, to establish the main result, showing that for any $E^{T}$, the period-1 conjectures $\{\varphi_{\hat{\mathrm{DS}},1}(\cdot,k):k\in A_{1}\cup B_{1}\}$ are nonempty and satisfy consistency is enough.
Fix an economy $E^{T}$. For each $k\in A_{1}\cup B_{1}$, let $\varphi_{0}(\emptyset,k)=\pazocal{M}_{T}(\emptyset,k)$ and recursively define for $n\geq 1$
$\varphi_{n}(\emptyset,k)=\left\{\overline{m}\in\pazocal{M}_{T}(\emptyset,k):\begin{array}[]{l}\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{n-1}}^{\overline{m}_{1}}),(\overline{m}_{t})_{t=2}^{T}|_{E_{2}^{T}(\overline{m}_{1})}\in\hat{\mathrm{DS}}_{T-1}(E_{2}^{T}(\overline{m}_{1}))\end{array}\right\}.$
We observe that for all $k\in A_{1}\cup B_{1}$, the following hold. First, $\varphi_{n+1}(\emptyset,k)\subseteq\varphi_{n}(\emptyset,k)$ because in each step we are increasing the continuation values of period-1 agents other than $k$. Formally, by definition we have that for all $k\in A_{1}\cup B_{1}$, $\varphi_{1}(\emptyset,k)\subseteq\varphi_{0}(\emptyset,k)$. Suppose we have shown that $\varphi_{n}(\emptyset,k^{\prime})\subseteq\varphi_{n-1}(\emptyset,k^{\prime})$ for all $k^{\prime}\in A_{1}\cup B_{1}$; we proceed to show that $\varphi_{n+1}\subseteq\varphi_{n}$. Fix $k\in A_{1}\cup B_{1}$ and $\overline{m}\in\varphi_{n+1}(\emptyset,k)$. We only need to show that if $\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{n}}^{\overline{m}_{1}})$, then $\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{n-1}}^{\overline{m}_{1}})$. To this end, let $k^{\prime}$ be such that $\overline{m}_{1}(k^{\prime})\neq k^{\prime}$. By construction, $k^{\prime}$ prefers $\overline{m}$ to the worst conjectured matching in $\varphi_{n}(\emptyset,k^{\prime})\subseteq\varphi_{n-1}(\emptyset,k^{\prime})$. Thus, $k^{\prime}$ also prefers $\overline{m}$ to the worst conjectured matching in $\varphi_{n-1}(\emptyset,k^{\prime})$. Thus, $\overline{m}\in\varphi_{n}(\emptyset,k)$. Second, because everything is finite $\varphi_{\infty}(\emptyset,k)=\lim_{n\rightarrow\infty}\varphi_{n}(\emptyset,k)$ is well-defined. Third, because $\{\overline{m}\in\pazocal{M}_{T}(\emptyset,k):(\overline{m}_{t})_{t=2}^{T}\in\hat{\mathrm{DS}}_{T-1}\left(E_{2}^{T}(\overline{m}_{1})\right)\}\subseteq\varphi_{n}(\emptyset,k)$ for all $n$ and is nonempty, then $\varphi_{\infty}(\emptyset,k)$ is nonempty.
@@
$\varphi_{\infty}(\emptyset,k)=\left\{\overline{m}\in\pazocal{M}_{T}(\emptyset,k):\begin{array}[]{l}\overline{m}_{1}\in\mathrm{S}(E_{1,\varphi_{\infty}}^{\overline{m}_{1}}),(\overline{m}_{s})_{s=2}^{T}|_{E_{2}^{T}(\overline{m}_{1})}\in\hat{\mathrm{DS}}_{T-1}(E_{2}^{T}(\overline{m}_{1}))\end{array}\right\},$ (B.2)
@@
    If $\overline{N}$ exists such that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\overline{N}}(\emptyset,k)=\varphi_{\overline{N}+1}(\emptyset,k)$, then for all $N\geq\overline{N}$ and all $k\in A_{1}\cup B_{1}$, it follows that $\varphi_{N}(\emptyset,k)=\varphi_{\overline{N}+1}(\emptyset,k)$. This follows from the definition of $\varphi_{n}$.
@@
    Because the sets $\varphi_{n}(\emptyset,k)$ are finite and so is the set of period-1 agents, $\overline{N}<\infty$ exists such that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\overline{N}}=\varphi_{\overline{N}+1}$. Toward a contradiction, suppose that for all $N\in\mathbb{N}$ $k_{N}\in A_{1}\cup B_{1}$ exists such that $\varphi_{N}(\emptyset,k_{N})\neq\varphi_{N+1}(\emptyset,k_{N})$. Since the set $A_{1}\cup B_{1}$ is finite, we can find an agent $k$ and a subsequence $(n^{k})$ such that $\varphi_{n^{k}}(\emptyset,k)\neq\varphi_{n^{k}+1}(\emptyset,k)$. Thus, $\varphi_{n^{k}+1}(\emptyset,k)\subsetneq\varphi_{n^{k}}(\emptyset,k)$. Since $\varphi_{n^{k}}(\emptyset,k)$ is finite, this means that eventually $\varphi_{n^{k}}(\emptyset,k)=\emptyset$, a contradiction. From now on, let $\overline{N}$ denote the smallest $n$ such that for all $k\in A_{1}\cup B_{1}$, $\varphi_{\overline{N}}(\emptyset,k)=\varphi_{\overline{N}+1}(\emptyset,k)\equiv\varphi_{\infty}(\emptyset,k)$.
@@
    Suppose $k\in A_{1}\cup B_{1}$ exists such that $\overline{m}\in\varphi_{\infty}(\emptyset,k)$, but $\overline{m}$ is not an element of the set on the right-hand side of Equation B.2. Then, without loss of generality, $a\in A_{1}\cup B_{1}\setminus\{k\}$ exists such that $\overline{m}_{1}(a)\neq a$ and $\min_{\tilde{m}\in\varphi_{\infty}(\emptyset,a)}U_{1}(a,\tilde{m})>u(a,\overline{m}_{1}(a)).$ This implies that $\overline{m}\notin\varphi_{\overline{N}+2}(\emptyset,k)$, contradicting that $\varphi_{\overline{N}+1}(\emptyset,k)=\varphi_{\infty}(\emptyset,k)$.
We conclude that for all $k\in A_{1}\cup B_{1}$, Equation B.2 holds and hence $\varphi_{\infty}=\varphi_{\hat{\mathrm{DS}},1}$. Thus, for all $k\in A_{1}\cup B_{1}$, $\varphi_{\hat{\mathrm{DS}}}(\emptyset,k)\neq\emptyset$.
We now show that $\varphi_{\hat{\mathrm{DS}},1}$ satisfy CC. To this end, construct a candidate $m^{\star}$ as described in Section 4 using $\varphi_{\hat{\mathrm{DS}},1}$ and the correspondence $\hat{\mathrm{DS}}_{T-1}$, which by the inductive hypothesis is nonempty. Fix a period-1 agent, $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)=k$. Note that $m^{\star}$ satisfies the following: First, the continuation matching $(m^{\star}_{t})_{t=2}^{T}\in\hat{\mathrm{DS}}_{T-1}(E^{T}_{2}(m^{\star}_{1}))$. Second, for all $k^{\prime}\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k^{\prime})\neq k^{\prime}$, we have that $k^{\prime}$ is matched to someone they prefer to the worst matching in $\varphi_{\hat{\mathrm{DS}},1}(\emptyset,k^{\prime})$. Finally, note that $m^{\star}_{1}$ has no blocking pairs $(a,b)$ such that $m^{\star}_{1}(a)\neq a,m^{\star}_{1}(b)\neq b$. Thus, $m^{\star}_{1}\in\mathrm{S}(E_{1,\varphi_{\hat{\mathrm{DS}}}}^{m^{\star}_{1}})$. Hence, $m^{\star}\in\varphi_{\hat{\mathrm{DS}},1}(\emptyset,k)$ so that CC holds.